OpenStudy (anonymous):

Write the first 6 terms of the sequence. a1=3 an=n^2+1

OpenStudy (anonymous):

a1=3; a2= 2^2+1=5 a3=3^2+1=10 a4=4^2+1=17 a5=5^2+1=26 a6=6^2+1=37 and so on...

OpenStudy (anonymous):

replace \(n\) by \(2\) then replace \(n\) by \(3\) etc

OpenStudy (anonymous):

@katherine.ok i don't understand why you did those steps? like why is a1=3 there?

OpenStudy (anonymous):

because you said a1= 3 yourself (perhaps it was because it was given in question)

OpenStudy (anonymous):

yes, it was given in the question

OpenStudy (anonymous):

@katherine.ok